I need help with these two problem please.

7 3/4 - 3 1/3 (7 and 3 over 4 (-minus) 3 and 1 over 2)

Hi Stanbon, thank you very much for your help. One thing, I don't understand the 7 and 3 over 4 -3 and 1 over 2 problem. I don't understand how you got the 14 over 2 at the end.

Rewrite the division problem before solving 1/4 divided 1/2
I need to know which one is correct:
1. 4/1 x 1/2

2. 4/1 x 2/1

3. 1/4 x 2/1

4. 1/4 x 1/2


Answer by checkley77(9435)

7 3/4 - 3 1/2
(4*7+3)/4-(2*3+1)/2
(28+3)/4-(6+1)/2
31/4-7/2
(2*31-7*4)/4*2
(62-28)/8
34/8=17/4 ANSWER.
--------------------------------
(1/4)/(1/2)
1/4*2/1=2/4=1/2
ANSWER=3)